# Your copula is a classifier in disguise: classification-based copula density estimation

### v1

@@ 3.2 Learning ratio copulas in practice
To obtain samples from $\mathcal{D}_{\tilde{q}}$, one could use the heuristic of recreating independent samples by dimension-wise shuffling the values of $\mathcal{D}_{\tilde{p}}$, but this would lead to a misrepresentation of the statistical properties. Instead, we can take advantage of our transformation to $\mathbb{R}^{d}$ in (5) and exploit the fact that ${\tilde{q}}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x}))=\mathcal{N}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\mathbf{0},\mathbf{I}_{d})$, which we can sample from ad nauseam. This satisfies our first desideratum, namely, access to a large amount of samples from $\tilde{q}$.
@@
We prove this statement in Appendix A.2. The result implies that under restrictions on the smallest width of the ellipsoid of $\tilde{p}$, the KL cannot be infinite for a fixed dimension $d$, and in fact scales linearly. As expected, if $\Sigma$ is close to nonsingular, the KL divergence can still be very large indeed; it is linear in $d$ but $\Sigma$ is likely to become more ill-conditioned (closer to nonsingularity) as $d$ grows. Consequently, we can rely on the empirical findings of Rhodes et al., (2020), who show classical DRE methods fail for KL differences of over 20 natural logarithms between the two densities. This result provides an informed way of deciding which DRE method to employ for a Gaussian copula based on the sample covariance of $\mathcal{D}_{\tilde{p}}$; whether to use a simple ratio copula model, empower it with a Gaussian-Guided ratio copula, or employ one of the many advanced DRE frameworks (Rhodes et al.,, 2020; Choi et al.,, 2021, 2022; Srivastava et al.,, 2023).
Diagnostics for overfitting
Other than using Lemma 3.2, an easily available diagnostic is to look at the loss during training. Since the copula naturally overlaps independent with dependent samples, we should always observe a positive loss. If it is too close to 0, then the classifier is overfitting and the two densities are too far apart (Rhodes et al.,, 2020). Furthermore, using a normalising constant estimator also provides a diagnostic of fit: copula densities are already normalised by definition, hence if correctly estimated, the normalising constant will be 1. If any of these symptoms arise, it signals we should employ a better DRE method for estimating the ratio copula. For instance, switching from a ratio copula to a Gaussian-Guided ratio copula.
@@ 4.1 Gaussian Copulas and QDA
In the case of a Gaussian copula, the density in the numerator is a correlated Gaussian and the denominator is an uncorrelated Gaussian, both centred at $\mathbf{0}$. From Equation 4, it then follows that the classification problem occurs between a multivariate Gaussian $\tilde{q}$ with identity covariance and a fully dependent Gaussian $\tilde{p}$. Thus, our aim becomes the classification of two Gaussians with different covariance structures. This is exactly the goal of Quadratic Discriminant Analysis (QDA) (Hastie et al.,, 2009), which fits a function $h^{k}(\mathbf{z};\mu_{1},\mu_{0},\Sigma_{1},\Sigma_{0}):\mathbb{R}^{d}\mapsto[0,1]$ to classify between $\mathcal{N}(.;\mu_{1},\Sigma_{1})$ with labels $y=1$ and $\mathcal{N}(.;\mu_{0},\Sigma_{0})$ with labels $y=0$.
@@
We provide a proof in Appendix A.3. This result opens up a new way of estimating Gaussian copulas by fitting a QDA classifier to $\Phi^{-1}\circ\mathcal{P}(\mathbf{x})$. Further, this result provides a new perspective on Gaussian copulas by linking them to the quadratic decision boundaries of QDA, characterising the complexity of Gaussian copulas.
@@ 4.2 In the case of a KDE copula
Next, we look at KDE copulas in terms of their implicit classification problems to establish their relative complexity. In particular, for a Gaussian KDE copula, we can obtain the following:
@@
The proof is included in Appendix A.4. Proposition 4.2 describes a KDE copula as an average of QDA decision boundaries over the $T_{p}$ observed samples. This results in a more complex decision boundary with more nonlinearity introduced linearly in the number of samples. However, due to the curse of dimensionality, an exponential number of samples will be required to maintain a constant level of complexity, thus elucidating KDE copulas’ difficulty in scaling to high dimensions.
@@ 4.3 For a simplified vine copula
Vine copulas are decompositions of copula densities using only bivariate copulas as building blocks and in simplified vine copulas, the bivariate copulas are assumed to have no conditioning (see Appendix B.2 and Czado et al., (2019) for an introduction). Here, we consider two cases, namely, using Gaussian copulas or Gaussian KDE copulas as those building blocks within simplified vine copula models. We give the statement here with a proof found in Appendix A.5.
@@
This result is instructive in describing the complexity of simplified vines through their use of multiple decision boundaries. They function much like hierarchical classifiers (examples of which are classification and regression trees, Loh, (2011), or random forests, Breiman, (2001)) by breaking the modelling task into smaller subparts on subgroups of variables. One can also draw a parallel with multilayered perceptrons that have parts of the model, such as their different layers, specialised in capturing specific features of the data relayed to them. Vine copulas instead have many separate models which capture different features, where components further up the decomposition have inputs with more conditioning - encoding the information extracted by previous classifiers. This can then be seen as a mixture of experts model, a common framework in machine learning (Yuksel et al.,, 2012; Jiang et al.,, 2024).
@@ 5 Experiments
GG+ratio $\mathbf{12.97\pm 0.67}$ $\mathbf{8.03\pm 0.05}$ $\mathbf{198.39\pm 19.72}$ $36.44\pm 0.51$
@@
Table 1: Average LL and W2 scores for the Digits and MNIST datasets with one standard deviation margins over 10 runs.
Section 4 described the limitations of popular copulas. To achieve better performance, designing more flexible copula models is essential. In this Section, we experimentally verify the validity of our proposed estimator. Implementation details are discussed in Appendix C, and our code is available at https://github.com/Huk-David/Ratio-Copula. We investigate the following:
Baselines:
We compare our GG ratio copula to commonly used models for copula estimation. Namely, the Gaussian copula, the TLL [^2] Transformation local likelihood (TLL), a class of nonparametric copula estimators using (5) with polynomials. constant, linear or quadratic copulas in the pyvinecopublib package of Vatter and Nagler, (2022) as well as a simplified vine copula using TLL constant bivariate copulas. Finally, we compare with the implicit generative copula (IGC) of Janke et al., (2021), as a comparison with an effective deep copula, which is only able to sample.
Metrics:
As a metric for success we choose the average copula log-likelihood (denoted LL), a well-used metric on its own, but in the context of copula models, once marginal densities have been estimated, only the copula LL is relevant to the LL of the full model. We also compare samples from each copula model using an empirical version of the Wasserstein-2 metric, as done in other probabilistic modelling contributions (e.g. Tagasovska et al., (2019)).
Sampling:
Our ratio copula offers an analytical density. Hence we use Hamiltonian Monte Carlo (Neal,, 2011), a gradient-informed MCMC sampler, to output samples from the estimated model.
Estimating 2D copulas:
In Figure 2 we fit parametric copulas based on 5000 samples, namely the Student-t, the Gumbell and a mixture of two Gaussian copulas with random covariances. We then test the model LL by evaluating 5000 held-out test samples, repeating this experiment 25 times with different seeds. We show the LL values for the Gaussian, TLL copulas, as well as two versions of our ratio copula: ratio copula (a multilayered perceptron (MLP)) and L+Ratio (a polynomial logistic classifier of order 5 with all interaction terms). The Gaussian copula performs the worst overall. The TLL and ratio copulas are similar for the Student-t, but the ratio copulas do better on the Gumbel and slightly worse on the Gaussian mixture. The estimator’s variance is higher than the TLL’s variance for the Gumbel, likely due to it being an Archimedean copula. Notably, the L+ratio copula marginally outperforms the MLP version, which we accredit to its simple structure acting as an implicit regularisation. Both ratio copulas perform consistently well, making them suitable for vine copulas.
Figure 2: Box plots showing the average LL across 25 runs on samples from three different copulas.
In Figure 3, we show estimates of toy examples of copulas generated from the dependence found in a 2D image when data is sampled according to the light of each pixel. The training size is one million, meant to emulate the big data setting. We show densities of the ratio copula, the TLL quad, and samples from the IGC. The ratio copula captures the sharp transitions of the densities accurately, while the TLL is much more diffuse. The IGC struggles to converge to a correct representation.
Figure 3: Bivariate copula models trained on the dependence in the light of pixels from 2D images.
@@
We model the copula density obtained with a 50/50 train/test split, from the digits dataset ($n=1797,d=64$) and the MNIST dataset ($n=60000,d=784$), both showing handwritten digits. This dataset assesses the ability of our model to pick up complex and multimodal dependencies in the data and is typically difficult for copula models. In Table 1, we show the results of the Gaussian copula, Gaussian-guided ratio copula, a simplified vine and the IGC. For digits, our method performs best in LL and sampling. For MNIST, we achieve a better LL, and are second in W2 to the IGC. We show samples of both digits and MNIST data from all methods in Appendix C.1. Notably, samples from the vine on MNIST are of worse quality, and IGC MNIST samples appear over-inflated or unfocused. Only the ratio copula captures the right thickness of the characters.
To conclude, we demonstrated that ratio copulas are extremely flexible models if given enough data (Figure 3), while in smaller data regimes, the L+ratio might be preferred. Thus, their use is warranted in vine constructions, broadening the class of non-parametric bivariate models. Further, though the application on Digits and MNIST, we demonstrate the ability of ratio copulas to scale to high dimensions with complex dependencies, thereby establishing ratio copulas as a useful non-parametric model.
@@ 6 Related Research
Our research elucidates connections in three ways. Firstly, this paper bridges copulas with DRE and the related Mutual information (MI) estimation. Secondly, in the taxonomy of implicit learning losses from Mohamed and Lakshminarayanan, (2016) applied to copulas, our work stands alone, not only through the use of a Bregman divergence, but also as the only deep copula using non-likelihood-based learning with the bespoke goal of density estimation. Lastly, we establish a connection with energy-based models (EBMs) and point out some existing implicit applications of copulas in the field.
@@
In Mohamed and Lakshminarayanan, (2016), three likelihood-free learning methods are discussed, with a shared goal of density estimation by comparison: [1] moment matching methods, [2] the maximum mean discrepancy (MMD), and [3] probabilistic classification (including Bregman and $f$-divergences). In the copula estimation community, likelihood-based training is prevalent. An exception is Hofert et al., (2021), who use a generative moment matching network, a purely generative model, to learn a deep copula by matching the moments of samples against observations - exemplifying an application of [1] to copulas. Similarly, Janke et al., (2021) use the MMD as a learning metric for a generative network to replicate observed copula samples - showcasing the use of [2]. Huk et al., (2023) use the MMD as a learning metric for when MLE-based training is unavailable but still obtain a likelihood model. Relatedly, Alquier et al., (2023) estimate parametric copulas with the MMD, instead motivated by its robustness properties. Finally, the current paper is the first copula estimation method implementing [3]: the use of probabilistic classification. Our work is unique in constructing a deep copula model trained using a likelihood-free objective while still obtaining a density estimator - in contrast to the contributions of Hofert et al., (2021) and Janke et al., (2021).
@@
EBMs (Ranzato et al.,, 2006) are a class of unormalised models giving unconstrained values to data on $\mathcal{X}$, to maximise the values given to observations from $p$ and minimise the values anywhere else. They are often trained with contrastive objectives (Hinton,, 2002), and in particular also employ probabilistic DRE (Gutmann and Hyvärinen,, 2012; Rhodes et al.,, 2020). They can be interpreted as an extreme version of ratio copulas, aimed at capturing dependence in data, not only discriminating against $\prod_{i}p_{i}$ but the whole space of $\mathcal{X}$. Conversely, our ratio copula model is an EBM trained on the ratio of $p/\prod_{i}p_{i}$. A notable use case of EBMs is modelling the latent space of variational autoencoders (VAE) Kingma, (2013), which is equipped with independent marginal densities. Xiao and Han, (2022) train an EBM with DRE to learn $p/\prod_{i}p_{i}$ for better performance, indirectly exemplifying the performance of ratio copulas! In fact, early classification DRE works advocate the use of $\prod_{i}p_{i}$ to learn $p$; Hastie et al., (2009) p.479 and Shi and Horvath, (2006) both unknowingly use a ratio copula model. Our paper thus consolidates their empirical work with a grounded derivation of this procedure, bridging the rich field of DRE with copula estimation.

### v2

@@ 3.2 Learning ratio copulas in practice
During training, all observations are labelled as dependent with $y=1$, whereas we have no independent samples from ${\tilde{q}}$. To obtain such samples, one could use the heuristic of recreating independent samples by dimension-wise shuffling the values of $\mathcal{D}_{\tilde{p}}$, but this would lead to a misrepresentation of the statistical properties. Instead, we can take advantage of our transformation to $\mathbb{R}^{d}$ in (5) and exploit the fact that ${\tilde{q}}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x}))=\mathcal{N}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x});\mathbf{0},\mathbf{I}_{d})$, which we can sample from ad nauseam. This satisfies our first desideratum, namely, access to a large amount of samples from $\tilde{q}$.
@@
We prove this statement in Appendix A.2. The result implies that under restrictions on the smallest width of the ellipsoid of $\tilde{p}$, the KL cannot be infinite for a fixed dimension $d$, and in fact scales linearly. As expected, if $\Sigma$ is close to nonsingular, the KL divergence can still be very large indeed; it is linear in $d$ but $\Sigma$ is likely to become more ill-conditioned (closer to nonsingularity) as $d$ grows. Consequently, we can rely on the empirical findings of Rhodes et al. [2020], who show classical DRE methods fail for KL differences of over 20 natural logarithms between the two densities. This result provides an informed way of deciding which DRE method to employ for a Gaussian copula based on the sample covariance of $\mathcal{D}_{\tilde{p}}$; whether to use a simple ratio copula model, empower it with a Gaussian-Guided ratio copula, or employ one of the many advanced DRE frameworks [Rhodes et al., 2020, Choi et al., 2021, 2022, Srivastava et al., 2023].
@@ 4 EQUIVALENCES BETWEEN CLASSIFIERS AND COPULAS
Further, in Figure 4 of Appendix A.7, we make use of the following identity
$c(\mathbf{P(\mathbf{x})})=\frac{h(\mathbf{x})}{1-h(\mathbf{x})}\iff h(\mathbf{x})=\frac{c(\mathbf{P(\mathbf{x})})}{1+c(\mathbf{P(\mathbf{x})})}$
to illustrate our theory by showing the implied classifiers $h(\mathbf{x})$ for six bivariate parametric copulas.
@@ 4.1 Gaussian Copulas and QDA
In the case of a Gaussian copula, the density in the numerator is a correlated Gaussian and the denominator is an uncorrelated Gaussian, both centred at $\mathbf{0}$. From Equation 4, it then follows that the classification problem occurs between a multivariate Gaussian $\tilde{q}$ with identity covariance and a fully dependent Gaussian $\tilde{p}$; our aim becomes the classification of two Gaussians with different covariance structures. This is exactly the goal of Quadratic Discriminant Analysis (QDA) [Hastie et al., 2009], which fits a function $h^{k}(\mathbf{z};\mu_{1},\mu_{0},\Sigma_{1},\Sigma_{0}):\mathbb{R}^{d}\mapsto[0,1]$ to classify between $\mathcal{N}(.;\mu_{1},\Sigma_{1})$ with labels $y=1$ and $\mathcal{N}(.;\mu_{0},\Sigma_{0})$ with labels $y=0$.
@@
We provide a proof in Appendix A.3. This result opens up a new way of estimating Gaussian copulas by fitting a QDA classifier to $\Phi^{-1}\circ\mathcal{P}(\mathbf{x})$. Further, this result provides a new perspective on Gaussian copulas by linking them to the quadratic decision boundaries of QDA, characterising the complexity of Gaussian copulas. We show an equivalent result for Student’s t copulas in Appendix A.4.
@@ 4.2 In the case of a KDE copula
Next, we look at KDE copulas regarding their implicit classification problems to establish their relative complexity. In particular, for a Gaussian KDE copula, we can obtain the following:
@@
$\hat{c}_{T_{p}}(\Phi^{-1}\circ\mathcal{P}(\mathbf{x}))=$ (11)
$\frac{\sum_{k=1}^{T_{p}}\prod_{i=1}^{d}\mathcal{N}(\Phi^{-1}\circ P_{i}(x_{i})-Z_{i}^{k}|0,b)}{T_{p}\cdot\mathcal{N}_{d}\left(\Phi^{-1}\circ\mathcal{P}(\mathbf{x})|\mathbf{0},\mathbf{I}_{d}\right)}.$ (12)
@@
The proof is included in Appendix A.5. This can be extended to other types of kernels for similar results. Proposition 4.2 describes a KDE copula as an average of QDA decision boundaries over the $T_{p}$ observed samples. This results in a more complex decision boundary with more nonlinearity introduced linearly in the number of samples. However, due to the curse of dimensionality, an exponential number of samples will be required to maintain a constant level of complexity, thus elucidating KDE copulas’ difficulty in scaling to high dimensions.
@@ 4.3 For a simplified vine copula
Vine copulas are decompositions of copula densities using only bivariate copulas as building blocks and in simplified vine copulas, the bivariate copulas are assumed to have no conditioning (see Appendix B.2 and Czado et al. [2019] for an introduction). Here, we consider two cases, namely, using Gaussian copulas or Gaussian KDE copulas as those building blocks within simplified vine copula models. We give the statement here with a proof found in Appendix A.6.
@@
This result is instructive in describing the complexity of simplified vines through their use of multiple decision boundaries. They function much like hierarchical classifiers (examples of which are classification and regression trees, Loh [2011], or random forests, Breiman [2001]) by breaking the modelling task into smaller subparts on subgroups of variables. One can also draw a parallel with multilayered perceptrons that have parts of the model, such as their different layers, specialised in capturing specific features of the data relayed to them. Vine copulas instead have many separate models which capture different features, where components further up the decomposition have inputs with more conditioning - encoding the information extracted by previous classifiers. This can then be seen as a mixture of experts model, a common framework in machine learning [Yuksel et al., 2012, Jiang et al., 2024].
@@ 5 EXPERIMENTS
Figure 2: Bivariate nonparametric copula models trained on the dependence in the light of pixels from 2D images. From left to right, columns show on $[0,1]^{2}$: the observed data, the LL of our ratio copula, samples from the IGC, and the LL of the TLL. Only the ratio copula manages to correctly capture the sharp transitions in the LLs.
Section 4 described the limitations of popular copulas. To achieve better performance, designing more flexible copula models is essential. In this Section, we experimentally verify the validity of our proposed estimator. Implementation details are discussed in Appendix C with additional experiments on alternate losses and samples produced by our ratio copula models. Our code is available at https://github.com/Huk-David/Ratio-Copula. We investigate the following:
Baselines:
We compare our ratio copula to commonly used models for copula estimation. Namely, the Gaussian copula, the TLL [^2] Transformation local likelihood (TLL), a class of nonparametric copula estimators using (5) with polynomials. constant, linear or quadratic copulas in the pyvinecopublib package of Vatter and Nagler [2022] as well as a simplified vine copula using TLL constant bivariate copulas. Finally, we compare with the implicit generative copula (IGC) of Janke et al. [2021], as a comparison with an effective deep copula, which is only able to sample.
Metrics:
As a metric for success we choose the average copula log-likelihood (denoted LL), a well-used metric on its own, but in the context of copula models, once marginal densities have been estimated, only the copula LL is relevant to the LL of the full model. We also compare samples from each copula model using an empirical version of the Wasserstein-2 metric, as done in other probabilistic modelling contributions (e.g. Tagasovska et al. [2019]).
Sampling:
Our ratio copula offers an analytical density. Hence we use Hamiltonian Monte Carlo [Neal, 2011], a gradient-informed MCMC sampler, to output samples from the estimated model.
Figure 3: Box plots showing the average LL across 25 fits of bivariate copula models based on samples from the Student’s t, Gumbel and Gaussian Mixture parametric copulas.
Estimating 2D copulas:
@@
GG+ratio ${12.97\pm 0.67}$ ${8.03\pm 0.05}$ ${198.39\pm 19.72}$ $36.44\pm 0.51$
GG+Ratio large $\mathbf{13.69\pm 0.77}$ $\mathbf{7.99\pm 0.07}$ $\mathbf{334.42\pm 45.91}$ $37.40\pm 0.79$
@@
Table 1: Average LL and W2 scores for the Digits and MNIST datasets with one standard deviation margins over 10 runs, the best scores are in bold. Our method (GG+Ratio) performs best and improves further with a larger network architecture for the classifier (GG+Ratio large), showcasing its adaptability.
In Figure 3 we fit parametric copulas based on 5000 samples, namely the Student-t, the Gumbell and a mixture of two Gaussian copulas with random covariances. We then test the model LL by evaluating 5000 held-out test samples, repeating this experiment 25 times with different seeds. We show the LL values for the Gaussian, TLL copulas, as well as two versions of our ratio copula: Ratio Copula (a multilayered perceptron) and L+Ratio (a polynomial logistic classifier of order 5 with all interaction terms). The Gaussian copula performs the worst overall. The TLL and our models are similar for the Student-t and on the Gumbel but perform slightly worse on the Gaussian mixture. This is a non-trivial achievement since current methods have been extensively optimized for low-dimensional data. Our estimator’s variance is higher than the TLL’s variance for the Gumbel, likely due to it being an Archimedean copula. Notably, the L+Ratio copula marginally outperforms the Ratio Copula version, which we accredit to its simple structure acting as an implicit regularisation. Both of our models perform consistently well, making them suitable for vine copulas.
In Figure 2, we show estimates of toy examples of bivariate copulas generated from the dependence found in monochromatic images when data is sampled according to the light of each pixel (more descriptions of this dataset can be found in C.2). The training size is one million, meant to emulate the big data setting. We show observed data in the first column, with the following columns showing the ratio copula density, samples from the IGC, and the TLL quad density, in that order. The ratio copula captures the sharp transitions of the densities accurately, while the TLL is much more diffuse. The IGC struggles to converge to a correct representation.
@@
We model the copula density obtained with a 50/50 train/test split, from the digits dataset ($n=1797,d=64$) and the MNIST dataset ($n=60000,d=784$), both showing handwritten digits. This dataset assesses the ability of our model to pick up complex and multimodal dependencies in the data and is typically difficult for copula models. In Table 1, we show the results of the Gaussian-guided ratio copula (GG+Ratio with a simple network and GG+Ratio large having a larger network architecture for the classifier, see Appendix C.3), a simplified vine, a Gaussian copula, and the IGC. Unlike the vine copula, whose building blocks scale exclusively with sample size (Corrolary 4.3), our method easily integrates more complex architectures as done with the GG+Ratio large.
For digits, our method performs best in LL and sampling. For MNIST, we achieve a better LL, and are second in W2 to the IGC. We show generated samples for both digits and MNIST from all methods in Appendix C.3. Notably, samples from the vine on MNIST are of worse quality, and IGC MNIST samples appear over-inflated or unfocused. Only the ratio copula captures the right thickness of the characters.
To conclude, we demonstrated that ratio copulas are extremely flexible models if given enough data (Figure 2), while in smaller data regimes, the L+ratio might be preferred. Thus, their use is warranted in vine constructions, broadening the class of non-parametric bivariate models. Further, though the application on Digits and MNIST, we demonstrate the ability of ratio copulas to scale to high dimensions with complex dependencies, thereby establishing ratio copulas as a useful non-parametric copula model.
@@ 6 RELATED RESEARCH
Our research elucidates connections in three ways. Firstly, this paper bridges copulas with DRE and the related Mutual information (MI) estimation. Secondly, in the taxonomy of implicit learning losses from Mohamed and Lakshminarayanan [2016] applied to copulas, our work stands alone, not only through the use of a Bregman divergence, but also as the only deep copula using non-likelihood-based learning with the bespoke goal of density estimation. Lastly, we establish a connection with energy-based models (EBMs) and point out some existing implicit applications of copulas in the field.
@@
In Mohamed and Lakshminarayanan [2016], three likelihood-free learning methods are discussed, with a shared goal of density estimation by comparison: [1] moment matching methods, [2] the maximum mean discrepancy (MMD), and [3] probabilistic classification (including Bregman and $f$-divergences). In the copula estimation community, likelihood-based training is prevalent. An exception is Hofert et al. [2021], who use a generative moment matching network, a purely generative model, to learn a deep copula by matching the moments of samples against observations - exemplifying an application of [1] to copulas. Similarly, Janke et al. [2021] use the MMD as a learning metric for a generative network to replicate observed copula samples - showcasing the use of [2]. Huk et al. [2023] use the MMD as a learning metric for when MLE-based training is unavailable but still obtain a likelihood model. Relatedly, Alquier et al. [2023] estimate parametric copulas with the MMD, instead motivated by its robustness properties. Finally, the current paper is the first copula estimation method implementing [3]: the use of probabilistic classification. Our work is unique in constructing a deep copula model trained using a likelihood-free objective while still obtaining a density estimator - in contrast to the contributions of Hofert et al. [2021] and Janke et al. [2021].
@@
EBMs [Ranzato et al., 2006] are a class of unormalised models giving unconstrained values to data on $\mathcal{X}$, to maximise the values given to observations from $p$ and minimise the values anywhere else. They are often trained with contrastive objectives [Hinton, 2002], and in particular also employ probabilistic DRE [Gutmann and Hyvärinen, 2012, Rhodes et al., 2020]. They can be interpreted as an extreme version of ratio copulas, aimed at capturing dependence in data, not only discriminating against $\prod_{i}p_{i}$ but the whole space of $\mathcal{X}$. Conversely, our ratio copula model is an EBM trained on the ratio of $p/\prod_{i}p_{i}$. A notable use case of EBMs is modelling the latent space of variational autoencoders (VAE) Kingma [2013], which is equipped with independent marginal densities. Xiao and Han [2022] train an EBM with DRE to learn $p/\prod_{i}p_{i}$ for better performance, indirectly exemplifying the performance of ratio copulas! In fact, early classification DRE works advocate the use of $\prod_{i}p_{i}$ to learn $p$; Hastie et al. [2009] p.479 and Shi and Horvath [2006] both unknowingly use a ratio copula model. Our paper thus consolidates their empirical work with a grounded derivation of this procedure, bridging the rich field of DRE with copula estimation.